MENTORING PAIRINGS

APPRENTICE Graduate students and e-mails

Name  @math.uchicago.edu
Kate Ponto  kponto
Alan Anders  aanders
Matt Booth  mbooth
Matt Holden  holden
Keerthi Madapusi (1-6)  keerthi
Deepam Patel (1-4)  dpatel
Boris Zbarsky (5-8)  bzbarsky

Organizer: Kate Ponto Undergraduate students, e-mails, and mentors.

Name (24)  e-mail  mentor(s)
Bargar, Emily  ebargar@uchicago.edu  Booth
Behr, Piotr S.  pbehr@uchicago.edu  Patel
Belyi, Alexander  abelyi@uchicago.edu  Holden
Chanyshev, Oleg  olegch@uchicago.edu  Holden
Chyr, Chingwei  chyr@uchicago.edu  Anders
Clarkson, Corrin  clarkson@uchicago.edu Anders
Gupta, Arpit  arpit@uchicago.edu Madapusi
Hansberger, Thomas J. thomash@uchicago.edu Madapusi
Janofsky, Eric  ebjan@uchicago.edu  Anders
Kreek, Robert A.  rakreek@uchicago.edu  Ponto
Mark, Alice  amark@uchicago.edu  Booth
Monson, Thomas  currence@uchicago.edu  Patel
Mussach, Jacqueline  jpeg@uchicago.edu  Holden
Noh, Hae Young  hynoh@uchicago.edu  Patel
Peysakhovich, Yana  yana@uchicago.edu  Holden
Picel, Jeffrey  jpicel@uchicago.edu  Ponto
Raghuram, Nikhil  nraghura@uchicago.edu Madapusi
Resnick, Rebecca  rebeccar@uchicago.edu Ponto
Salvatore, James jsal13@uchicago.edu Booth
Shirzadi, Cyrus  cyrus@uchicago.edu Madapusi
Shron, Maxwell  mshron@uchicago.edu  Booth
Wang, Lisa  lywang@uchicago.edu  Ponto
Yu, David  dryu@uchicago.edu  Patel
Zhong, Wayne (Weizhen) waynez@uchicago.edu Anders

SPECIAL SIT-INS

Norton, Emily brntnrtn@uchicago.edu Ponto
Rezvi, Sara  arsinoe@uchicago.edu Ponto

2 MENTORING PAIRINGS

YSP Graduate students and e-mails

Name  @math.uchicago.edu
Mohammed Abouzaid (1, 5-8)  mabouzai
Catalin Carstea  carstea
David Diamondstone  ded
Jim Fowler (1-2, 5-8)  fowler
Aaron Marcus  amarcus
Emma Smith (1-2, 5-8)  emmas
Robert Young  rjyoung

Organizer: Mohammed Abouzaid Undergraduate students, e-mails, and mentors.

Name (22)  email  Program  mentor(s)
Altschul, Sam  salt@uchicago.edu  11-12  Marcus
Baraki, Semere G.  sgb@uchicago.edu  7-8  Carstea
Bender, Gabriel  gbender@uchicago.edu  11-12  Young
Bruce, Matthew  mbruce@uchicago.edu  7-8  Carstea
Chariker, Christopher L. chariker@uchicago.edu  9-10 Smith/Diamondstone
Constantin, Sarah  sconstan@Princeton.EDU  7-8  Smith/Diamondstone
Flood, Stephen  sflood@uchicago.edu  7-8  Diamondstone
Ford, Nicolas  nford@uchicago.edu  9-10  Fowler/Marcus
Gardner, Daniel  dbgardner@uchicago.edu 11-12  Fowler/Marcus
Gaster, Gabriel  eulerian@uchicago.edu  9-10  Fowler/Marcus
Gunther, Joseph  jgunther@uchicago.edu  9-10  Diamondstone
Henry, Shawn  shawnh@uchicago.edu  11-12  Young
Kalikman, Philip  plk@uchicago.edu  11-12 Smith/Diamondstone
Lin, Zhiwei Calvin  linzw@uchicago.edu  9-10  Young
Meyer, Jeff  jsmeyer@uchicago.edu  7-8  Smith/Diamondstone
Nieder, Jonathan  jrnieder@uchicago.edu  9-10  Fowler/Marcus
Raskin, Samuel  sraskin@uchicago.edu  9-10  Fowler/Marcus
Ruggeri, Edward  ruggeri@uchicago.edu  7-8  Marcus
Schonert, Jeff  jeffrey1@uchicago.edu  11-12  Carstea
Sheldon, Abigail  sheldona@uchicago.edu  7-8  Smith/Diamondstone
Tan, Weiwei  cora7@uchicago.edu  9-10  Young
Taylor, Brian  bgtaylor@uchicago.edu  7-8  Carstea

MENTORING PAIRINGS 3

SESAME Graduate students and e-mails

Name  @math.uchicago.edu
Ian Biringer  biringer
Shawn Drenning  shawn
Tullia Dymarz (5-8)  tullia
Masoud Kamgarpour  mkamgarp
Justin Noel  justin
Andy Putman  andyp
Saravanan Thiyagarajan  avan
Dani Zarnescu  zarnescu

Organizer: Mohammed Abouzaid Undergraduate students, e-mails, and mentors.

Name (26)  e-mail  mentor(s)
Balady, Stephen  sbalady@uchicago.edu  Biringer
Bartlett, Padraic James  padraic@uchicago.edu  Putman
Casarotto, Chad  chadmbol@uchicago.edu  Noel
Davis, Trevor L.  trevorld@uchicago.edu  Zarnescu
Dickinson, Dan  dand@uchicago.edu  Biringer
Doss, Charles  cdoss@uchicago.edu  Biringer
Eisenberg, Andy  aeisen@uchicago.edu  Kamgarpour
Gonwa, Matthew  gonwa@uchicago.edu  Thiyagarajan
Herin, Sayer  sayerh@uchicago.edu  Thiyagarajan
Herschlag, Gregory J.  Gtzu@uchicago.edu  Drenning
Hudec, Drew  dhudec@uchicago.edu  Putman
Immerman, Daniel  immerman@uchicago.edu Thiyagarajan
Kidd, Milan  mkidd@uchicago.edu  Putman
Kiersz, Andrew  akiersz@uchicago.edu  Noel
Levi, Asa  asa@uchicago.edu  Biringer
McNeur, Joshua C.  jmcneur@uchicago.edu  Noel
Miller, Nilsen  nilsen1@uchicago.edu  Drenning
O’Meara, Matthew  momeara@uchicago.edu  Zarnescu
Prag, Robert  rprag@uchicago.edu  Zarnescu
Raskin, Sarah  sarahr@uchicago.edu  Kamgarpour
Redelsheimer, Katrina kredelsheimer@uchicago.edu  Putman
Steiner, Andrew M.  steinera@uchicago.edu  Noel
Sternberg, John  jds3@uchicago.edu  Kamgarpour
Thvedt, Michael  cynicalk@uchicago.edu  Drenning
Waliji, Muhammad  mhwaliji@uchicago.edu  Zarnescu
Weiner, Dan  dbw@uchicago.edu  Kamgarpour